OpenStudy (anonymous):

Many sweepstakes contests have an elimination round. In the elimination round, half of the entrants are chosen at random to go on to the final round. Then one person is chosen as the winner. What is the theoretical probability that a person who enters a contest with 10, 000 entrants will be the winner?

OpenStudy (anonymous):

@agent0smith @AccessDenied @Ashleyisakitty @adrynicoleb @Awesome781 @Ash90 @beccaboo333 @bbbbbbbbbbbbbbbbbb @bbcream14 @Bedubbs @bbcream14 @berlingots

OpenStudy (kropot72):

the probability that an entrant will be chosen to go into the final round is 1/2. The probability of an entrant in the final round will be the winner is 1/5,000. Therefore the probability that a person who enters a contest with 10, 000 entrants will be the winner is given by: \[\frac{1}{2}\times\frac{1}{5000}=you\ can\ calculate\]

OpenStudy (anonymous):

@Compassionate @chocodropa7 @Cjsmall14 @cheergurl_99 @CaseyCarns @chrisdbest @divu.mkr @Derrickhitch @dominick6541 @D.udeM.eetsT.houghts @DoeBoiiMontana @Deaner @ememlove @edwardp @eighthourlunch @eighthourlunch @einie @Elegant_Beautyy @ellemoraes

OpenStudy (agent0smith):

If you enter a contest with 10,000 people and only 1 wins, it doesn't matter how many elimination rounds they do...your chance is 1 in 10,000 regardless.